Please attempt both.1. Perform the marquis de Laplace process in both possible ways (remember, you choose the lead vector) on the basis 3 -2 -1 t
Answer #1

ANSWER

0 0 Given 604[:] [:]}={u, ugy **01011 (-2)*(-1)-2)x(1) = 2-3 = -1 0 so, the sets is linearly independent and hence normal bas*- 444 3 = 14 5 7/5 13-14 15-14 t+7/5) -5+7 So, we can check (-2,1). () - 2 + 2 = 0 59-41:] . So tha vezuinol orthogonal basi3 (-) + (-2)/- 1 - 1 / +-) (-1) 425/3=114 +(-3)??? = 60°/-231+5)*8-1.3(-1+3)+1-3 (5+3) with respect to istrow = 42)(6)-3()+31二 -H+15 - 2 +1+9 3 3 (3) 1-11-11) 又十 colit dle ele 29 乡 介 5 | 1 4-4 1 35-12 四4 =15 g 23 59 v. - 29/4 -V4 Q911V = -1 12913 ,M3= 4 231 VB =U3-U3V VINI V) - La V2 V 벗 (1,1)-213) 13 (31,3) (31,-) () ) (끝) t 29 It 구 -6-1- 3 -- HET t8 4+ +9100 tho + 5그 19키 It 23] 3- 10 블 be XSI 71 + 25X-11 tx it 네 5 구 25X23 txit 3 10 125 497 1411 - 귀 0-725 497 It + 275 497 - 497so, Required orthogona basis for R3 ૧૧h ૬), ૧૧,